# On-the-fly Surrogation for Complex Nonlinear Dynamics*

### v2

@@ IV Examples
In this section, we present numerical examples to show the capabilities of the proposed on-the-fly surrogation approach. For this purpose, we provide two different examples. In the first example, we consider the Van der Pol oscillator, and our objective is to provide a clear and intuitive demonstration of the proposed method. In the second example, we consider an interconnection of mass-spring-damper (MSD) systems to show the scalability of the proposed approach. The full implementation of these examples is available at: https://gitlab.com/Javi-Olucha/cdc25-code-repo
@@ IV-A Van der Pol oscillator
$\Sigma_{\mathrm{vdp}}\coloneq\bigl{\{}\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -x_{1}-\eta x_{2}(1-x_{1}^{2})\end{bmatrix}+\begin{bmatrix}0\\ x_{1}u\end{bmatrix},$ (31)
where $\eta>0$ is a system parameter. Moreover, let $\Sigma_{\mathrm{vdp}}^{\tilde{\eta}}$ denote the Van der Pol oscillator with the parameter $\eta$ set at a fixed value $\tilde{\eta}>0$. Furthermore, we consider the infinite-horizon optimal control law given by $u=-x_{1}x_{2}$ for the performance objective $J(x,u)=x_{2}^{2}+u^{2}$ proposed by [27] that drives the system to the origin.
@@ IV-A1 Reconstruction of the true nonlinear system from the linear dynamics
First, we demonstrate the best achievable performance from the methodology itself with the reconstruction of the global dynamics of the Van der Pol oscillator $\Sigma_{\mathrm{vdp}}^{0.5}$ based on perfect reconstruction of the underlying local linearizations. In this case, as the origin of $\Sigma_{\mathrm{vdp}}^{0.5}$ is an equilibrium, we can use the results in Corollary 2. Then, with the optimal control law $u$, we execute a CT simulation of $\Sigma_{\mathrm{vdp}}^{0.5}$ for each of the initial conditions $(-2,-2)$, $(-2,2)$, $(2,-2)$ and $(2,2)$.
Figure 1: State responses of the CT simulation of the Van der Pol oscillator with $\eta=0.5$, where ( ) is the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$ and (   ) is the reconstruction of $\Sigma_{\mathrm{vdp}}^{0.5}$ using the results in Corollary 2 with a perfect reconstruction of the underlying local linearizations.
For the simulations, we use the Matlab in-built variable-step solver ode45 with the default parameters and step size, and simulate for 14 seconds. Next, we replicate the simulations using the same settings, but this time we use (12) instead of the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$. For the integral in (12) we use the Matlab in-built function integral with the default tolerances, which computes a numerical approximation of the integral based on the global adaptive quadrature approach from [28]. For these simulations, ode45 requested an average of 21300 evaluations of the Jacobian of (31) with an average computation time of 0.16 seconds [^1] On a 64 bit laptop PC with 2.10 GHz 8-core CPU and 64 GB RAM. The simulation results, shown in Fig. 1, illustrate that the solutions of $\Sigma_{\mathrm{vdp}}^{0.5}$ and its reconstruction with Corollary 2 are identical, with an average root-mean-square error (RMSE) of $3.3\times 10^{-16}$, supporting the results in Subsection III-A.
@@ IV-A2 On-the-fly surrogation of the Van der Pol oscillator
Second, we find a surrogate of $\Sigma_{\mathrm{vdp}}^{0.5}$ using the proposed on-the-fly surrogation method. For this, we define the following set of $N=9$ points:
$\small\Theta_{\mathrm{vdp}}^{1}=\left\{\!\begin{array}[]{cc}(-2,-2,-4),&(-2,2,4),\\ (2,-2,4),&(2,2,-4),\\ (-1.8,-1.8,-3.24),&(-1.8,1.8,3.24),\\ (1.8,-1.8,3.24),&(1.8,1.8,-3.24),\\ (0,0,0)\end{array}\!\right\}.$ (32)
@@
Next, we construct the set $\mathcal{M}_{\mathrm{vdp}}^{1}$ by linearizing $\Sigma_{\mathrm{vdp}}^{0.5}$ at each point in $\Theta_{\mathrm{vdp}}^{1}$ and define the set $\mathcal{D}_{\mathrm{vdp}}^{1}=\{\Theta_{\mathrm{vdp}}^{1},\mathcal{M}_{\mathrm{vdp}}^{1}\}$. Then, we construct the interpolant $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$ with Algorithm 1, with the RBF $\phi$ given in (27) and the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\beta_{3}x_{2}+\beta_{4}u$. The width parameter $c$ is determined by minimizing (30) with Bayesian optimization [29], since (30) is a non-smooth function. Note that other optimization methods suitable for non-smooth objectives can also be employed. Finally, we construct the surrogate representations $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$ and $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$ using $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$ and Algorithm 2 with two distinct numerical integration schemes. These two schemes are integral from Matlab, used for $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$, and the composite Simpson’s 3/8 rule (CS38) [30, Chapter 5.1] with 6 equidistant intervals, used for $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$.
@@
Figure 2: State responses of the CT simulation of the Van der Pol oscillator with $\eta=0.5$, where ( ) is obtained with the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$, ( ) with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$ and (   ) with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$.
Now, we compare the surrogates w.r.t. $\Sigma_{\mathrm{vdp}}^{0.5}$ by CT simulations, where the simulation and numerical integration settings for integral are set as described in the previous experiment IV-A1. For the simulations with $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{int}}$, ode45 requested an average of 20400 evaluations of $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$, leading to an average computation time and RMSE of 0.655 seconds and 0.1, respectively. For the simulations with $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}\mathrm{cs38}}$, ode45 requested an average of 816 evaluations of $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$, leading to an average computation time and RMSE of 0.028 seconds and 0.1, respectively. The results are shown in Fig. 2, where we observe that with only $N=9$ distinct linearization points, the obtained surrogates can approximate the behaviour of $\Sigma_{\mathrm{vdp}}^{0.5}$. Moreover, the use of CS38 drastically reduces the computation time, while the effects in numerical accuracy are negligible.
@@ IV-A3 On-the-fly surrogation of the Van der Pol oscillator for a range of system parameter values
Third, we consider that the system parameter $\eta$ in (31) lies within the range $\eta\in[0.3,0.6]$. To capture the system dynamics for the different values of $\eta$, we define the set $\mathcal{D}_{\mathrm{vdp}}^{2}$, which contains $N=68$ system linearizations contained within the admissible spaces:
$\mathcal{X}_{\mathrm{vdp}}$ =$=\{\underline{x}_{i}=-2.2\leq x_{i}\leq\overline{x}_{i}=2.2\}_{i=1}^{2},$ (33)
$\mathcal{U}_{\mathrm{vdp}}$ =$=\{\underline{x}_{1}\underline{x}_{2}\leq u=-x_{1}x_{2}\leq-\overline{x}_{1}\overline{x}_{2}\},$
$\mathcal{H}_{\mathrm{vdp}}$ =$=\{\underline{\eta}=0.3\leq p\leq\overline{\eta}=0.6\}.$
Now, we construct the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ and compare in simulation with the true nonlinear system, using the same settings for the interpolant and the simulations described in the second experiment IV-A2, and CS38 with 6 equidistant intervals. In addition, for these simulations we set $\eta$ to different fixed values contained in (33) that are not included in $\Theta_{\mathrm{vdp}}^{2}$. In this case, ode45 also requested a maximum of 816 evaluations of the interpolant, leading to an average computation time and RMSE of 0.045 seconds and 0.784, respectively. These results are displayed in Fig. 3.
Figure 3: State responses of the CT simulation of the Van der Pol oscillator, where ( ) and ( ) are obtained with the true nonlinear systems ${\Sigma}_{\mathrm{vdp}}^{0.33}$ and ${\Sigma}_{\mathrm{vdp}}^{0.47}$, respectively, and (   ) and (   ) are obtained with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ evaluated at $\eta=0.33$ and $\eta=0.47$, respectively.
@@ IV-B Interconnection of MSD systems
$\Sigma_{\mathrm{msd}}\coloneq\left\{\begin{aligned} &m\ddot{x}_{1}=-F_{1}-F_{1,2};\\ &m\ddot{x}_{i}=-F_{i}-F_{i,\,i-1}-F_{i,\,i+1};\\ &m\ddot{x}_{\mathrm{Nm}}=-F_{\mathrm{Nm}}-F_{\mathrm{Nm},\,\mathrm{Nm}-1}+u,\end{aligned}\right.$ (34)
@@
for $12\leq i\leq\mathrm{Nm}-$ 1, where $F_{i,\,j}$ and $F_{i}$ are the forces applied to the $i$-th mass by the adjacent masses and from the connection to the rigid wall, respectively, given by
$\begin{split}F_{i,j}&=k_{1}(x_{i}\!-\!x_{j})\!+\!b_{1}(x_{i}\!-\!x_{j})\!+\!(x_{i}\!-\!x_{j})^{3}\!+\!b_{2}(\dot{x}_{i}\!-\!\dot{x}_{j})^{3};\\ F_{i}&=k_{1}x_{i}+b_{1}\dot{x}_{i},\end{split}$
and the system parameters are chosen as the number of masses $\mathrm{Nm}=5$, $m=1$ kg, $k_{1}=0.5$ N/m, $b_{1}=1$ Ns/m, $b_{2}=2$ Ns/m, and the external force $u$ applied to the last mass is in Newtons. Further, we define the admissible spaces
$\mathcal{X}_{\mathrm{msd}}$ =$=\left\{\!\begin{array}[]{c}\{-2.2\leq x_{i}\leq=2.2\}_{i=1}^{5},\\ \{-1.5\leq x_{i}\leq=1.5\}_{i=6}^{10}\end{array}\!\right\};$ (35)
$\mathcal{U}_{\mathrm{msd}}$ =$=\{-1.5\leq u\leq=1.5\},$
@@
and we construct the data set $\mathcal{D}_{\mathrm{msd}}$ with $N=100$ system linearizations at randomly drawn distinct points uniformly distributed within (35). Note that, as the respective state-order and input dimension of $\Sigma_{\mathrm{msd}}$ are $n_{\mathrm{x}}=10$ and $n_{\mathrm{u}}=1$, a regular grid containing only the vertices of the hyper-rectangle described by (35) requires $20482^{n_{\mathrm{x}}+n_{\mathrm{u}}}\!=\!$ 2048 system linearizations.
@@
Similar as in Subsection IV-A2, we construct the interpolant $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, where $\mathcal{D}_{\mathrm{msd}}$, the RBF $\phi$ in (27), the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\dots\beta_{11}x_{10}+\beta_{12}u$, and the width parameter $c$ obtained by minimizing (30) with Bayesian optimization are the inputs to Algorithm 1. Then, we formulate the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ according to Algorithm 2 using $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ and CS38 with 6 equidistant intervals.
@@
Next, we test $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ by executing one thousand CT simulations, where the states corresponding to the mass velocities, i.e., $\{x_{i}\}_{i=6}^{10}$ are initialized at zero, and the states corresponding to the mass positions, i.e., $\{x_{i}\}_{i=1}^{5}$ are initialized at random values drawn within the admissible space defined in (35) with a uniform distribution. Moreover, the external force $u$ is set to $u(t)=0.7\sin(2\pi t)$, and each simulation is executed for 8 seconds using ode45 with the default parameters and step size. For these simulations, ode45 requested an average of 2352 evaluations of $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, leading to an average computation time of 0.15 seconds. In Fig. 5
@@
Figure 6: Expected probability density distribution of the RMSE over one thousand simulations of $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. $\Sigma_{\mathrm{msd}}$, denoted as $e(\cdot,\cdot)$, where the expected probability density distribution is computed based on a normal kernel function and ( ) indicates the mean of the error.
we display the trajectory of the last mass obtained from some simulation results, where it is observed that the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ approximates true nonlinear system $\Sigma_{\mathrm{msd}}$ with remarkable accuracy. Furthermore, we compute the RMSE for each state of the simulated $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. the simulated $\Sigma_{\mathrm{msd}}$. Then, we compute the expected probability density distribution of the RMSE data based on a normal kernel function [31], and besides some outliers with error values of up to $6\cdot 10^{-3}$, most of the error across all simulations is below $10^{-3}$, as shown in Fig. 6.

### v3

@@ IV Examples
In this section, we present two numerical examples to show the capabilities of the proposed on-the-fly surrogation approach. The first example considers the Van der Pol oscillator to provide a clear and intuitive demonstration of the method. The second example involves an interconnection of mass-spring-damper (MSD) systems to demonstrate scalability. The full implementation is available at: https://gitlab.com/Javi-Olucha/cdc25-code-repo.
@@ IV-A Van der Pol oscillator
$\Sigma_{\mathrm{vdp}}\coloneq\left\{\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -x_{1}-\eta x_{2}(1-x_{1}^{2})\end{bmatrix}+\begin{bmatrix}0\\ x_{1}u\end{bmatrix}\right$., (30)
where $\eta>0$ is a system parameter. Moreover, let $\Sigma_{\mathrm{vdp}}^{{\eta}}$ denote the Van der Pol oscillator with the parameter $\eta$ fixed at ${\eta}>0$. Furthermore, we consider the infinite-horizon optimal control law given by $u=-x_{1}x_{2}$ for the performance objective $J(x,u)=x_{2}^{2}+u^{2}$ proposed by [22] that drives the system to the origin.
@@ IV-A1 Reconstruction of the true nonlinear system from the linear dynamics
First, we numerically demonstrate Theorem 1 with the reconstruction of the global dynamics of the Van der Pol oscillator $\Sigma_{\mathrm{vdp}}^{0.5}$ based on perfect reconstruction of the underlying local dynamics [^1] For complex models, obtaining the local linearized dynamics at any point in the admissible space is too expensive in practise, and only a limited set of local linearizations is available.. In this case, as the origin of $\Sigma_{\mathrm{vdp}}^{0.5}$ is an equilibrium, we use Corollary 2. Then, we execute a CT simulation of $\Sigma_{\mathrm{vdp}}^{0.5}$ for each of the initial conditions (2,2)(-2,-2), (2,2)(-2,2), (2,2)(2,-2) and $(2,2)$. For the simulations, we use the Matlab built-in variable-step solver ode45 with the default parameters and step size, and simulate for 14 seconds. Next, we replicate the simulations using the same settings, but this time we use (12) instead of the true nonlinear system $\Sigma_{\mathrm{vdp}}^{0.5}$. The integral in (12) is solved numerically using the composite Simpson’s 3/8 rule (CS38) [23, Chapter 5.1] with 6 equidistant intervals. For these simulations, ode45 requested an average of 852 evaluations of the Jacobian of (30) with an average computation time of 0.0035 seconds [^2] On a laptop with an i7-13850HX (2.10 GHz) CPU and 64 GB RAM.. The simulation results, shown in Fig. 1, illustrate that the solutions of $\Sigma_{\mathrm{vdp}}^{0.5}$ and its reconstruction with Corollary 2 are identical, with an average root-mean-square error (RMSE) of $2.38\times 10^{-16}$, supporting the results in Subsection III-A.
Figure 1: Numerical illustration of Theorem 1 using complete knowledge of the linearization operator. The state responses of the CT simulation of $\Sigma_{\mathrm{vdp}}^{0.5}$ are shown, where ( ) is obtained with the true system representation and (   ) with its reconstruction using Corollary 2.
@@ IV-A2 On-the-fly surrogation of the Van der Pol oscillator
We next learn surrogate models of $\Sigma_{\mathrm{vdp}}^{0.5}$ using the proposed on-the-fly surrogation method. For this, we use the sets $\Theta^{1}_{\mathrm{vdp}}\subset\Theta^{2}_{\mathrm{vdp}}\subset\Theta^{3}_{\mathrm{vdp}}$ defined in Tab. I, which contain 9, 15 and 25 observation points, respectively.
TABLE I: Definition of different sets with observation points used to obtain different surrogate models of $\Sigma_{\mathrm{vdp}}^{0.5}$
$\Theta^{1}_{\mathrm{vdp}}=\{(x_{1},x_{2})\mid x_{1}\in\{-2,0,2\},\,x_{2}\in\{-2,0,2\}\}$
$\Theta^{2}_{\mathrm{vdp}}=\{(x_{1},x_{2})\mid x_{1}\in\{-2,0,2\},\,x_{2}\in\mathbb{Z},\;-2\leq x_{2}\leq 2\}$
$\Theta^{3}_{\mathrm{vdp}}=\{(x_{1},x_{2})\mid x_{1},x_{2}\in\mathbb{Z},\;-2\leq x_{1}\leq 2,\;-2\leq x_{2}\leq 2\}\}$
@@
From these, we construct the dictionaries $\mathcal{D}_{\mathrm{vdp}}^{i}=\{\Theta_{\mathrm{vdp}}^{i},\mathcal{M}_{\mathrm{vdp}}^{i}\}$ with $i=1,2,3$, where the sets $\mathcal{M}_{\mathrm{vdp}}^{i}$ are obtained by linearizing $\Sigma_{\mathrm{vdp}}^{0.5}$ at the points in $\Theta_{\mathrm{vdp}}^{i}$. The interpolants $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{1}}$, $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{2}}$ and $\mathcal{I}_{\mathrm{vdp}}^{\mathcal{D}_{3}}$ are then computed using Algorithm 1 with the RBF $\phi$ given in (26) and the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\beta_{3}x_{2}$. The width parameter $c$ is determined by minimizing (29) with Bayesian optimization [24], given that (29) is non-smooth. Other non-smooth optimization schemes could also be employed. Next, the OTFS surrogate representations $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$, $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ and $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{3}}$ are built with Algorithm 2, the CS38 numerical integration rule with 6 intervals and the corresponding interpolants.
For comparison, we compute the sixth- and tenth-order Koopman surrogates $\hat{\Sigma}_{\mathrm{vdp}}^{K_{6}}$ and $\hat{\Sigma}_{\mathrm{vdp}}^{K_{10}}$, using the data-driven method of [25]. Both models are trained with the PyKoopman library [26] on the trajectories of Fig. 1 sampled with a time interval of $0.01$ seconds. The observables $g(x)=[1,x_{1},x_{2},x_{1}^{2},x_{1}x_{2},x_{2}^{2}]^{\top}$ are used for $\hat{\Sigma}_{\mathrm{vdp}}^{K_{6}}$, while $\hat{\Sigma}_{\mathrm{vdp}}^{K_{10}}$ employs a third-order extension of $g(x)$.
@@
Figure 2: State responses of the CT simulation of $\Sigma_{\mathrm{vdp}}^{0.5}$, where ( ) is obtained with the true system, (   ), (   ) and (   ) with the respective OTFS surrogates $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$, $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ and $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{3}}$, and (     ) with the data-driven Koopman model $\hat{\Sigma}_{\mathrm{vdp}}^{K}$.
We compare the surrogates with $\Sigma_{\mathrm{vdp}}^{0.5}$ by CT simulations under the same settings as in the previous experiment IV-A1. The results, displayed in Fig. 2, show that all the obtained surrogates approximate the behaviour of $\Sigma_{\mathrm{vdp}}^{0.5}$. Remarkably, $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$ -trained with only 9 system snapshots- already outperforms the accuracy of the Koopman surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{K}$. Furthermore, as predicted by Theorem 5, the accuracy of the OTFS models improves systematically with richer data dictionaries. Average RMSEs and runtimes are given in Tab. II.
TABLE II: RMSE between the true system and the surrogate models and runtimes for the simulation results of Fig.  2.
Surrogate model Average RMSE Average runtime (s)
$\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{1}}$ 0.1187 0.0466
$\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ 0.0743 0.0477
$\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{3}}$ 0.008 0.0493
$\hat{\Sigma}_{\mathrm{vdp}}^{K_{6}}$ 0.2404 0.0054
$\hat{\Sigma}_{\mathrm{vdp}}^{K_{10}}$ 0.0272 0.0062
@@ IV-A3 On-the-fly surrogation of the Van der Pol oscillator for a range of system parameter values
Third, we consider that the system parameter $\eta$ in (30) lies within the range $\eta\in[0.3,0.6]$. To capture the system dynamics for the different values of $\eta$, we define the observation set
$\Theta_{\mathrm{vdp}}^{4}=\{(x_{1},x_{2},\eta)\ \mid\ x_{1},x_{2}\in\mathbb{Z},-2\leq x_{1}\leq 2,$ (31)
[…]
and obtain the data dictionary $\mathcal{D}_{\mathrm{vdp}}^{4}$. Now, we obtain the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{4}}$ and compare in simulation with the true nonlinear system, using the same settings for the interpolant and the simulations described in the second experiment IV-A2, and CS38 with 6 equidistant intervals. In addition, for these simulations we fix $\eta$ to different values contained in (31) that are not included in $\Theta_{\mathrm{vdp}}^{2}$. In this case, ode45 requested a maximum of 978 evaluations of the interpolant, leading to an average runtime and RMSE of 0.048 seconds and 0.106, respectively.
Figure 3: State responses of the CT simulation of the Van der Pol oscillator, where ( ) and ( ) are obtained with the true nonlinear systems ${\Sigma}_{\mathrm{vdp}}^{0.35}$ and ${\Sigma}_{\mathrm{vdp}}^{0.47}$, respectively, and (   ) and (   ) are obtained with the surrogate $\hat{\Sigma}_{\mathrm{vdp}}^{\mathcal{D}^{2}}$ evaluated at $\eta=0.35$ and $\eta=0.47$, respectively.
These results are displayed in Fig. 3.
@@ IV-B Interconnection of MSD systems
$\Sigma_{\mathrm{msd}}\coloneq\left\{\begin{aligned} &m\ddot{x}_{1}=-F_{1}-F_{1,2},\\ &m\ddot{x}_{i}=-F_{i}-F_{i,\,i-1}-F_{i,\,i+1},\\ &m\ddot{x}_{n_{\mathrm{m}}}=-F_{n_{\mathrm{m}}}-F_{n_{\mathrm{m}},\,n_{\mathrm{m}}-1}+u,\end{aligned}\right$. (32)
@@
for $2\leq i\leq n_{\mathrm{m}}-1$, where $F_{i,\,j}$ and $F_{i}$ are the forces applied to the $i$-th mass by the adjacent masses and from the connection to the rigid wall, respectively, given by
$\begin{split}F_{i,j}&=k_{1}(x_{i}\!-\!x_{j})\!+\!b_{1}(x_{i}\!-\!x_{j})\!+\!(x_{i}\!-\!x_{j})^{3}\!+\!b_{2}(\dot{x}_{i}\!-\!\dot{x}_{j})^{3},\\ F_{i}&=k_{1}x_{i}+b_{1}\dot{x}_{i},\end{split}$
@@
and the system parameters are chosen as the number of masses $n_{\mathrm{m}}=5$, $m=1$ kg, $k_{1}=0.5$ N/m, $b_{1}=1$ Ns/m, $b_{2}=2$ Ns/m, and the external force $u$ applied to the last mass is in Newtons. Further, we define the admissible space $\mathcal{Z}_{\mathrm{msd}}:=\mathcal{X}_{\mathrm{msd}}\times\mathcal{U}_{\mathrm{msd}}$, where $\mathcal{X}_{\mathrm{msd}}:=\{\{-2.2\leq x_{i}\leq 2.2\}_{i=1}^{5},\{-1.5\leq x_{i}\leq 1.5\}_{i=6}^{10}\}$, $\mathcal{U}_{\mathrm{msd}}=\{-1.5\leq u\leq 1.5\}$, and we construct the data set $\mathcal{D}_{\mathrm{msd}}$ with $N=100$ distinct system linearizations at randomly drawn points uniformly distributed within $\mathcal{Z}_{\mathrm{msd}}$. In contrast, as the respective state-order and input dimension of $\Sigma_{\mathrm{msd}}$ are $n_{\mathrm{x}}=10$ and $n_{\mathrm{u}}=1$, a regular grid containing only the vertices of the hyper-rectangle described by $\mathcal{Z}_{\mathrm{msd}}$ requires $2^{n_{\mathrm{x}}+n_{\mathrm{u}}}\!=\!2048$ system snapshots.
@@
Similar as in Subsection IV-A2, we construct the interpolant $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, where $\mathcal{D}_{\mathrm{msd}}$, the RBF $\phi$ in (26), the polynomial basis $\mathcal{I}_{\mathrm{poly}}(x,u)=\beta_{1}+\beta_{2}x_{1}+\dots\beta_{11}x_{10}+\beta_{12}u$, and the width parameter $c$ obtained by minimizing (29) with Bayesian optimization are the inputs to Algorithm 1. Then, we assemble the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ according to Algorithm 2 using $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ and CS38 with 6 equidistant intervals.
@@
Next, we test $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ by executing one thousand CT simulations, where the states corresponding to the mass velocities, i.e., $\{x_{i}\}_{i=6}^{10}$, are initialized at zero, and the states corresponding to the mass positions, i.e., $\{x_{i}\}_{i=1}^{5}$, are initialized at random values drawn within $\mathcal{Z}_{\mathrm{msd}}$ with a uniform distribution. Moreover, the external force $u$ is set to $u(t)=0.7\sin(2\pi t)$, and each simulation is executed for 8 seconds using ode45 with the default parameters and step size. For these simulations, ode45 requested an average of 2352 evaluations of $\mathcal{I}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$, leading to an average computation time of 0.18 seconds. In Fig. 5
@@
Figure 6: Estimated pdf of the RMSE over one thousand simulations of $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. $\Sigma_{\mathrm{msd}}$, denoted as $e(\cdot,\cdot)$, where the pdf is computed based on a normal kernel function and ( ) indicates the mean of the error.
we display the trajectory of the last mass obtained from some simulation results, where it is observed that the surrogate $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ approximates true nonlinear system $\Sigma_{\mathrm{msd}}$ with remarkable accuracy. Furthermore, we compute the RMSE for each state of the simulated $\hat{\Sigma}_{\mathrm{msd}}^{\mathcal{D}_{\mathrm{msd}}}$ w.r.t. the simulated $\Sigma_{\mathrm{msd}}$. Then, we compute the estimated probability density function (pdf) of the RMSE data based on a normal kernel function [27], and besides some outliers with error values of up to $6\cdot 10^{-3}$, most of the error across all simulations is below $10^{-3}$, as shown in Fig. 6.